OpenStudy (jenniferfournierr):

Confirm that f and g are inverses by showing that f(g(x)) = x and g(f(x)) = x. f of x equals two divided by x and g of x equals two divided by x

ILovePuppiesLol (ilovepuppieslol):

@.Sam. could probably help

ILovePuppiesLol (ilovepuppieslol):

@Abhisar this one

OpenStudy (zzr0ck3r):

\(f(x)=\dfrac{2}{x}\) \(g(x)=\dfrac{2}{x}\) \(f(g(x))=\dfrac{2}{f(x)}=\dfrac{2}{\frac{2}{x}}=2*\frac{x}{2}=?\) Do something similar for \(g(f(x))\)